# Forecasting and Manipulating the Forecasts of Others

### 12 Mar 2026 v1

@@ 5.1 Model and embedding
Define the aggregate primitive-noise demand kernel $\mathcal{D}_{\Sigma}(t,s):=\sum_{j}\mathcal{D}^{j}(t,s)$. Since $E^{V}E^{0,\top}=0$, all price information comes through drift-based inference:
@@
where $\widetilde{C}^{0}$ is the market maker’s unresolved drift kernel under the projection decomposition (A.12). The execution cost $\mathbb{E}[\int D_{t}^{i}P_{t}\,dt]$ is quadratic in $D_{t}^{i}$ with coefficient $\lambda(t)$—the endogenous $G^{i}_{DD}$.
@@ 5.2 Equilibrium characterization
$\underbrace{\sigma_{V}E^{V}\,\partial_{u}\bigl(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0}\bigr)(u)}_{H^{i,\mathrm{val}}(t,u):\;\text{information rent density}}\;=\;\underbrace{\int_{t}^{T}\delta P_{\tau}\,D^{i}(\tau,u)\,d\tau}_{H^{i,\mathrm{imp}}(t,u):\;\text{impact cost density}},$ (5.3)
where $\delta P_{\tau}$ is the price perturbation for an impulse at $t$. Differentiating both sides in $t$ and solving for $D^{i}(t,u)$ yields a closed backward ODE:
$\lambda(t)\,D^{i}(t,u)=-\frac{d}{dt}H^{i,\mathrm{val}}(t,u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}(\tau,u)\,d\tau,$ (5.4)
with terminal condition $D^{i}(T,u)=\lambda(T)^{-1}H^{i,\mathrm{val}}(T,u)$.
@@
This requires that aggregate order flow remains informative about the fundamental at every date, the economically relevant case since temporary uninformativeness would present an arbitrage opportunity. The condition is self-reinforcing: positive $\lambda$ induces finite trading rates, which maintain the drift informativeness that sustains $\lambda>0$. It plays the role of $G^{i}_{DD}\succ 0$ (Lemma 3.4), ensuring strict convexity of the best-response problem on the domain of the fixed-point operator.
@@ 5.3 Two channels of belief manipulation
    $\mathcal{V}^{i,\mathrm{price}}(t,u)=\lambda(t)\int_{t}^{T}D^{i}(\tau,u)\,d\tau$: instantaneous impact times the integrated future policy kernel. This is the standard Kyle channel—the trader internalizes how each demand impulse moves the price.
@@
    $\mathcal{V}^{i,\mathrm{opp}}(t,u)=\int_{t}^{T}\delta P_{\tau}^{\mathrm{ind}}\,D^{i}(\tau,u)\,d\tau$: opponent-mediated impact, arising because opponents also observe order flow and adjust their trading, which feeds back into prices.
@@ 6 Conclusion
A single coordinate change—conditioning on primitive Brownian shocks—reduces Nash equilibrium in continuous-time LQG games with endogenous signals to a deterministic fixed point in two-time kernels, with no truncation and no large-population limit. The information wedge $\mathcal{V}^{i}_{t}$—with mean component $\bar{\mathcal{V}}^{i}(t)$ and kernel component $\mathcal{V}^{i}(t,r)$—prices bilateral belief manipulation and vanishes when signals are exogenous. The two-player example shows that nearly all welfare gains from mandatory disclosure come from eliminating this channel; the Kyle–Back embedding shows the noise-state technique extends beyond quadratic costs and identifies a second channel—opponent inference through shared order flow—that applies immediately to asymmetric precisions and continuous private learning.
@@
The main open question is whether the deterministic-kernel class is without loss of generality. A natural program proceeds in three steps: (i) a Blackwell reduction showing that under quadratic costs the best response depends on the posterior $\pi_{t}^{i}:=\mathcal{L}(W_{[0,t]}\mid\mathcal{F}_{t}^{i})$ only through its mean and covariance; (ii) a belief-of-belief collapse in which player $i$’s estimate of $\pi_{t}^{k}$ factors through a deterministic operator—the posterior analogue of the kernel maps $(F^{i},\widetilde{X}^{i})$; (iii) a contraction argument on the variance functional $V(\mathbf{D}):=\sum_{i}\mathbb{E}\int\!\!\int\|D^{i}(t,s,\pi_{t}^{i})-\mathbb{E}[D^{i}(t,s,\pi_{t}^{i})]\|^{2}\,ds\,dt$, which vanishes exactly in the deterministic-kernel class and which iterated best responses are expected to contract. Beyond this, stochastic coefficients (Remark C.2), graph-structured interactions, and global uniqueness via the dissipative structure of $F^{i}$ are natural extensions.
@@ Appendix A Filtering
Throughout, formal calculations identify the kernel equations for $F^{i}$; we then define $F^{i}$ as the unique solution of those equations, with uniqueness following from uniqueness of $L^{2}$-projection. The pipeline proceeds in three steps. A.1 derives the dynamics of $t\mapsto\widehat{W}_{t}^{i}(u)$ from a characterizing identity for conditional expectations, using an ansatz with undetermined coefficients matched against exponential test martingales. A.2 specializes the abstract dynamics to the physical observation model, identifying $\widetilde{X}^{i}(t,u)$ as the filtering gain via Itô isometry, establishing the algebraic identity $\widetilde{X}^{i}=X(I-\Pi^{i})-X*F^{i}$, and recording the contraction identity $\int_{0}^{t}X(t,s)\,\widetilde{X}^{i}(t,s)^{\top}\,ds=\Sigma_{XX}^{i}(t)$. A.3 defines $F^{i}$ as the primitive filtering object through its forward evolution equation, derives the explicit three-term formula, and verifies uniqueness.
@@ A.1 Dynamics of the noise-state
Proposition A.1 (Volterra closure for $\widehat{W}$).
The map $u\mapsto\widehat{W}_{t}^{i}(u)$ is a Volterra process: there exists a deterministic kernel $F_{t}^{i}(\cdot,\cdot)$ such that
@@ A.2 Specialization to the physical observation model
which corresponds to setting $C(t,s)=\sqrt{P^{i}(t)}\,X(t,s)$ in the general framework of Proposition A.2.
@@
Substituting $C(t,s)=\sqrt{P^{i}(t)}\,X(t,s)$ into Proposition A.2 and using $dI_{t}^{i}=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$, the dynamics (A.2) become
$d_{t}\widehat{W}_{t}^{i}(u)=\Big(\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,X(t,s)^{\top}\,ds\Big)\sqrt{P^{i}(t)}\,dI_{t}^{i}\,+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}}\,dI_{t}^{i}.$ (A.9)
@@
$d_{t}\!\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du,\qquad 0\leq u<t\leq T.$ (A.10)
To verify the gain coefficient, recall that $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}(t,s)\,dW_{s}$ and that the innovation (A.17) has martingale component $E^{i}\,dW_{t}$ and absolutely continuous component $\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}(t,s)\,dW_{s}\,dt$. The gain in (A.10) equals the conditional cross-covariance density $\operatorname{Cov}(dW_{u},\,\sqrt{P^{i}(t)}(X_{t}{-}\widehat{X}_{t}^{i})\,dt\mid\mathcal{F}_{t}^{i})/dt$; by Itô isometry applied to $\int_{0}^{t}\widetilde{X}^{i}(t,s)\,dW_{s}$, this picks out $\widetilde{X}^{i}(t,u)$. Thus $\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}$ is the Kalman gain for updating the $u$-th noise increment: it maps observation innovations at time $t$ into revisions of the noise estimate at the earlier time $u$.
@@
$\widetilde{X}^{i}(t,u)\;=\;\frac{\operatorname{Cov}\!\big(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i}\big)}{du}.$
@@
$\operatorname{Cov}\!\big(X_{t},\,W_{u}\mid\mathcal{F}_{t}^{i}\big)\;=\;\int_{0}^{u}\widetilde{X}^{i}(t,s)\,ds\;=\;\int_{0}^{t}X(t,s)\,\partial_{s}\mathsf{C}_{t}^{i}(s,u)\,ds,$ (A.11)
where the second equality follows from $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}X(t,s)\,d\widetilde{W}_{t}^{i}(s)$ and Itô isometry against $\widetilde{W}_{t}^{i}(u)$.
@@
Recall the induced estimated-state kernel $\widehat{X}^{\,i}(t,s):=X(t,s)\Pi^{i}+\int_{0}^{t}X(t,z)\,F_{t}^{i}(z,s)\,dz$. Since $\widetilde{X}^{i}(t,u):=X(t,u)-\widehat{X}^{\,i}(t,u)$,
$\widetilde{X}^{i}(t,u)=X(t,u)(I-\Pi^{i})-\int_{0}^{t}X(t,z)\,F_{t}^{i}(z,u)\,dz.$ (A.12)
@@
$\Sigma_{XX}^{i}(t)=\int_{0}^{t}X(t,s)\,\widetilde{X}^{i}(t,s)^{\top}\,ds,$ (A.13)
which follows from Itô isometry applied to $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}(t,s)\,dW_{s}$.
@@
Identity (A.13) says that the $d\times d$ state uncertainty $\Sigma_{XX}^{i}(t)$ is obtained by “projecting” the unresolved kernel $\widetilde{X}^{i}(t,s)$ back through the state impulse response $X(t,s)$. In the single-player case with no endogenous drift, this reduces to the classical relation between the Kalman gain and the error covariance. In the multi-player setting the identity persists because the forward kernel environment is deterministic.
@@ A.3 The filtering kernel $F^{i}$
$\partial_{t}F_{t}^{i}(u,s)$ $=\widetilde{X}^{i}(t,u)^{\top}P^{i}(t)\,\widetilde{X}^{i}(t,s),\qquad u,s<t,$ (A.14)
$F_{t}^{i}(u,t)$ $=\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}(t,u).$ (A.15)
@@
$F_{t}^{i}(u,s)=\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr.$ (A.16)
@@ A.3.1 Derivation
$d_{u}\widehat{W}_{t}^{i}(u)=E^{i,\top}\,dI_{u}^{i}+\Bigg(\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}\Bigg)\,du,$
@@
The first piece is the martingale component; the second is absolutely continuous. Substituting the Volterra forms and using $\widetilde{X}^{i}(t,r):=X(t,r)-\widehat{X}^{\,i}(t,r)$ (the definition of $\widetilde{X}^{i}$ and (A.13)):
$dI_{t}^{i}=E^{i}\,dW_{t}+\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}(t,r)\,dW_{r}\,dt.$ (A.17)
@@
$E^{i,\top}\,dI_{u}^{i}$ $=\underbrace{E^{i,\top}E^{i}\,dW_{u}}_{=\;\Pi^{i}\,dW_{u}}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\int_{0}^{u}\widetilde{X}^{i}(u,r)\,dW_{r}\,du.$ (A.18)
The first piece contributes the $\Pi^{i}\,dW_{u}$ in the density representation. The second is absolutely continuous in $u$ with primitive-noise density $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)$ at source-noise index $s\leq u$.
@@
Substituting (A.17) into $\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$:
$\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$ $=\underbrace{\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\,dW_{s}}_{\text{(a) martingale}}$
$\quad+\;\underbrace{\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}P^{i}(s)\int_{0}^{s}\widetilde{X}^{i}(s,r)\,dW_{r}\,ds}_{\text{(b) abs.\ continuous}}.$
Term (a) contributes $\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$ at source-noise index $s\in(u,t]$. Applying stochastic Fubini to term (b) and reading off the coefficient of $dW_{r}$:
$\text{(b)}=\int_{0}^{t}\Bigg(\int_{\max(u,r)}^{t}\widetilde{X}^{i}(s,u)^{\top}P^{i}(s)\,\widetilde{X}^{i}(s,r)\,ds\Bigg)\,dW_{r}.$
@@
Collecting the coefficient of $dW_{s}\,du$ from all three contributions and noting that causality ($\widetilde{X}^{i}(s,u)=0$ for $u>s$) automatically zeroes the irrelevant indicators:
$F_{t}^{i}(u,s)=\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr.$ (A.19)
@@
  • $\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$: the unresolved response at observation time $s$, projected through the observation channel. Nonzero only when $u\leq s$ (causal in the observation direction).
  • $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)$: the drift component of the innovation at time $u$, which reveals information about earlier noise increments at index $s\leq u$. Nonzero only when $s\leq u$ (causal in the noise direction).
  • $\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr$: accumulated indirect inference—the joint information about the $u$-th and $s$-th increments extracted from observations at times $r\in[\max(u,s),\,t]$.
@@
Evaluating (A.19) at $s=t$ (using $\widetilde{X}^{i}(u,t)=0$ for $u<t$):
$F_{t}^{i}(u,t)=\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}(t,u),$
@@
$\partial_{t}F_{t}^{i}(u,s)=\widetilde{X}^{i}(t,u)^{\top}P^{i}(t)\,\widetilde{X}^{i}(t,s),$
@@
The formula (A.19) is explicit in the gain kernel $\widetilde{X}^{i}$: once $\widetilde{X}^{i}$ is computed from (A.12), the filtering kernel $F^{i}$ requires no further fixed-point iteration. In the numerical scheme of Section 4, the integral $\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr$ is accumulated forward alongside the $\widetilde{X}^{i}$-computation, and the two boundary terms are set at each new observation time.

### 18 Mar 2026 v2

@@ 5.1 Model and embedding
Define the aggregate primitive-noise demand kernel $\mathcal{D}_{\Sigma,t}(s):=\sum_{j}\mathcal{D}^{j}(t,s)$. Since $E^{V}E^{0,\top}=0$, all price information comes through drift-based inference:
@@
where $\widetilde{C}^{0}$ is the market maker’s unresolved drift kernel under the projection decomposition (A.12).
Remark 5.1 (Quadratic structure of the profit).
Fix opponents’ strategy maps. Since the market maker’s filtering gains are deterministic, the price decomposes as $P_{t}=P_{t}^{-i}+\int_{0}^{t}\ell(t,s)\,D_{s}^{i}\,ds$, where $P_{t}^{-i}$ is a linear process independent of $D^{i}$ and $\ell(t,s)$ is a deterministic kernel. Symmetrically, the future execution cost of a unit impulse at time $t$ is $\int_{t}^{T}D_{\tau}^{i}\,\delta P_{\tau}\,d\tau$, where $\delta P_{\tau}$ propagates forward through the deviation system with deterministic coefficients. The profit $\pi^{i}=\mathbb{E}[\int_{0}^{T}D_{t}^{i}(V_{T}-P_{t})\,dt]$ is therefore concave quadratic in $D^{i}$, when $\lambda(t)>0$ provides strict concavity.
@@ 5.2 Equilibrium characterization
$\underbrace{\sigma_{V}E^{V}\,\partial_{u}\bigl(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0}\bigr)(u)}_{H^{i,\mathrm{val}}_{t}(u):\;\text{information rent density}}\;=\;\underbrace{\int_{t}^{T}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau}_{H^{i,\mathrm{imp}}_{t}(u):\;\text{impact cost density}},$ (5.3)
where $\delta P_{\tau}$ is the price perturbation for an impulse at $t$. Differentiating both sides in $t$ and solving for $D^{i}_{t}(u)$ yields a closed backward ODE:
$\lambda(t)\,D^{i}_{t}(u)=-\frac{d}{dt}H^{i,\mathrm{val}}_{t}(u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau,$ (5.4)
with terminal condition $D^{i}_{T}(u)=\lambda(T)^{-1}H^{i,\mathrm{val}}_{T}(u)$.
@@
This requires that aggregate order flow remains informative about the fundamental at every date, the economically relevant case since temporary uninformativeness would present an arbitrage opportunity. The condition is self-reinforcing: positive $\lambda$ induces finite trading rates, which maintain the drift informativeness that sustains $\lambda>0$. It plays the role of $G^{DD,i}\succ 0$ (Lemma 3.4), ensuring strict convexity of the best-response problem on the domain of the fixed-point operator.
@@ 5.3 Two channels of belief manipulation
    $\mathcal{V}^{i,\mathrm{price}}(t,u)=\lambda(t)\int_{t}^{T}D^{i}_{\tau}(u)\,d\tau$: instantaneous impact times the integrated future policy kernel. This is the standard Kyle channel—the trader internalizes how each demand impulse moves the price.
@@
    $\mathcal{V}^{i,\mathrm{opp}}(t,u)=\int_{t}^{T}\delta P_{\tau}^{\mathrm{ind}}\,D^{i}_{\tau}(u)\,d\tau$: opponent-mediated impact, arising because opponents also observe order flow and adjust their trading, which feeds back into prices.
@@ 6 Conclusion
Conditioning on primitive Brownian shocks collapses the Townsend belief hierarchy onto deterministic impulse-response maps, yielding the first exact equilibrium characterization of finite-player continuous-time LQG games with endogenous signals. The information wedge $\mathcal{V}^{i}_{t}$ prices bilateral belief manipulation and vanishes when signals are exogenous. The two-player example shows that nearly all welfare gains from disclosure come from eliminating this channel; the Kyle–Back embedding shows the technique extends beyond quadratic costs, identifying a second channel through opponent inference via shared order flow that applies to asymmetric precisions and continuous private learning.
@@
The main open question is whether the deterministic impulse-response class is without loss of generality. A natural program proceeds in three steps: (i) a Blackwell reduction showing that under quadratic costs the best response depends on the posterior $\pi_{t}^{i}:=\mathcal{L}(W_{[0,t]}\mid\mathcal{F}_{t}^{i})$ only through its mean and covariance; (ii) a belief-of-belief collapse in which player $i$’s estimate of $\pi_{t}^{k}$ factors through a deterministic operator—the posterior analogue of the kernel maps $(F^{i},\widetilde{X}^{i})$; (iii) a contraction argument on the variance functional $V(\mathbf{D}):=\sum_{i}\mathbb{E}\int\!\!\int\|D^{i}(t,s,\pi_{t}^{i})-\mathbb{E}[D^{i}(t,s,\pi_{t}^{i})]\|^{2}\,ds\,dt$, which vanishes exactly in the deterministic impulse-response class and which iterated best responses are expected to contract. Beyond this, stochastic coefficients (Remark C.2), graph-structured interactions, and global uniqueness via the dissipative structure of $F^{i}$ are natural extensions.
@@ Appendix A Filtering
Throughout, formal calculations identify the kernel equations for $F^{i}$; we then define $F^{i}$ as the unique solution of those equations, with uniqueness following from uniqueness of $L^{2}$-projection. The pipeline proceeds in three steps. A.1 derives the dynamics of $t\mapsto\widehat{W}_{t}^{i}(u)$ from a characterizing identity for conditional expectations, using an ansatz with undetermined coefficients matched against exponential test martingales. A.2 specializes the abstract dynamics to the physical observation model, identifying $\widetilde{X}^{i}_{t}(u)$ as the filtering gain via Itô isometry, establishing the algebraic identity $\widetilde{X}^{i}=X(I-\Pi^{i})-X*F^{i}$, and recording the contraction identity $\int_{0}^{t}X_{t}(s)\,\widetilde{X}^{i}_{t}(s)^{\top}\,ds=\Sigma^{XX,i}(t)$. A.3 defines $F^{i}$ as the primitive filtering object through its forward evolution equation, derives the explicit three-term formula, and verifies uniqueness.
@@ A.1 Dynamics of the noise-state
Proposition A.1 (Linear closure for $\widehat{W}$).
The map $u\mapsto\widehat{W}_{t}^{i}(u)$ is a linear process: there exists a deterministic indirect-response map $F_{t}^{i}(\cdot,\cdot)$ such that
@@ A.2 Specialization to the physical observation model
which corresponds to setting $C(t,s)=\sqrt{P^{i}(t)}\,X_{t}(s)$ in the general framework of Proposition A.2.
@@
Substituting $C(t,s)=\sqrt{P^{i}(t)}\,X_{t}(s)$ into Proposition A.2 and using $dI_{t}^{i}=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$, the dynamics (A.2) become
$d_{t}\widehat{W}_{t}^{i}(u)=\Big(\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,X_{t}(s)^{\top}\,ds\Big)\sqrt{P^{i}(t)}\,dI_{t}^{i}\,+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}}\,dI_{t}^{i}.$ (A.9)
@@
$d_{t}\!\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du,\qquad 0\leq u<t\leq T.$ (A.10)
To verify the gain coefficient, recall that $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$ and that the innovation (A.17) has martingale component $E^{i}\,dW_{t}$ and absolutely continuous component $\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}\,dt$. The gain in (A.10) equals the conditional cross-covariance density $\operatorname{Cov}(dW_{u},\,\sqrt{P^{i}(t)}(X_{t}{-}\widehat{X}_{t}^{i})\,dt\mid\mathcal{F}_{t}^{i})/dt$; by Itô isometry applied to $\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$, this picks out $\widetilde{X}^{i}_{t}(u)$. Thus $\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}$ is the Kalman gain for updating the $u$-th noise increment: it maps observation innovations at time $t$ into revisions of the noise estimate at the earlier time $u$.
@@
$\widetilde{X}^{i}_{t}(u)\;=\;\frac{\operatorname{Cov}\!\big(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i}\big)}{du}.$
@@
$\operatorname{Cov}\!\big(X_{t},\,W_{u}\mid\mathcal{F}_{t}^{i}\big)\;=\;\int_{0}^{u}\widetilde{X}^{i}_{t}(s)\,ds\;=\;\int_{0}^{t}X_{t}(s)\,\partial_{s}\mathsf{C}_{t}^{i}(s,u)\,ds,$ (A.11)
where the second equality follows from $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}X_{t}(s)\,d\widetilde{W}_{t}^{i}(s)$ and Itô isometry against $\widetilde{W}_{t}^{i}(u)$.
@@
Recall the induced estimated-state kernel $\widehat{X}^{i}_{t}(s):=X_{t}(s)\Pi^{i}+\int_{0}^{t}X_{t}(z)\,F_{t}^{i}(z,s)\,dz$. Since $\widetilde{X}^{i}_{t}(u):=X_{t}(u)-\widehat{X}^{i}_{t}(u)$,
$\widetilde{X}^{i}_{t}(u)=X_{t}(u)(I-\Pi^{i})-\int_{0}^{t}X_{t}(z)\,F_{t}^{i}(z,u)\,dz.$ (A.12)
@@
$\Sigma^{XX,i}(t)=\int_{0}^{t}X_{t}(s)\,\widetilde{X}^{i}_{t}(s)^{\top}\,ds,$ (A.13)
which follows from Itô isometry applied to $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$.
@@
Identity (A.13) says that the $d\times d$ state uncertainty $\Sigma^{XX,i}(t)$ is obtained by “projecting” the unresolved kernel $\widetilde{X}^{i}_{t}(s)$ back through the state impulse response $X_{t}(s)$. In the single-player case with no endogenous drift, this reduces to the classical relation between the Kalman gain and the error covariance. In the multi-player setting the identity persists because the forward kernel environment is deterministic.
@@ A.3 The filtering kernel $F^{i}$
$\partial_{t}F_{t}^{i}(u,s)$ $=\widetilde{X}^{i}_{t}(u)^{\top}P^{i}(t)\,\widetilde{X}^{i}_{t}(s),\qquad u,s<t,$ (A.14)
$F_{t}^{i}(u,t)$ $=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}_{t}(u).$ (A.15)
@@
$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.16)
@@ A.3.1 Derivation
$d_{u}\widehat{W}_{t}^{i}(u)=E^{i,\top}\,dI_{u}^{i}+\Bigg(\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}\Bigg)\,du,$
@@
The first piece is the martingale component; the second is absolutely continuous. Substituting the linear forms and using $\widetilde{X}^{i}_{t}(r):=X_{t}(r)-\widehat{X}^{i}_{t}(r)$ (the definition of $\widetilde{X}^{i}$ and (A.13)):
$dI_{t}^{i}=E^{i}\,dW_{t}+\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}_{t}(r)\,dW_{r}\,dt.$ (A.17)
@@
$E^{i,\top}\,dI_{u}^{i}$ $=\underbrace{E^{i,\top}E^{i}\,dW_{u}}_{=\;\Pi^{i}\,dW_{u}}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\int_{0}^{u}\widetilde{X}^{i}_{u}(r)\,dW_{r}\,du.$ (A.18)
The first piece contributes the $\Pi^{i}\,dW_{u}$ in the density representation. The second is absolutely continuous in $u$ with primitive-noise density $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)$ at source-noise index $s\leq u$.
@@
Substituting (A.17) into $\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$:
$\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$ $=\underbrace{\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\,dW_{s}}_{\text{(a) martingale}}$
$\quad+\;\underbrace{\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}P^{i}(s)\int_{0}^{s}\widetilde{X}^{i}_{s}(r)\,dW_{r}\,ds}_{\text{(b) abs.\ continuous}}.$
Term (a) contributes $\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$ at source-noise index $s\in(u,t]$. Applying stochastic Fubini to term (b) and reading off the coefficient of $dW_{r}$:
$\text{(b)}=\int_{0}^{t}\Bigg(\int_{\max(u,r)}^{t}\widetilde{X}^{i}_{s}(u)^{\top}P^{i}(s)\,\widetilde{X}^{i}_{s}(r)\,ds\Bigg)\,dW_{r}.$
@@
Collecting the coefficient of $dW_{s}\,du$ from all three contributions and noting that causality ($\widetilde{X}^{i}_{s}(u)=0$ for $u>s$) automatically zeroes the irrelevant indicators:
$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.19)
@@
  • $\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$: the unresolved response at observation time $s$, projected through the observation channel. Nonzero only when $u\leq s$ (causal in the observation direction).
  • $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)$: the drift component of the innovation at time $u$, which reveals information about earlier noise increments at index $s\leq u$. Nonzero only when $s\leq u$ (causal in the noise direction).
  • $\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr$: accumulated indirect inference—the joint information about the $u$-th and $s$-th increments extracted from observations at times $r\in[\max(u,s),\,t]$.
@@
Evaluating (A.19) at $s=t$ (using $\widetilde{X}^{i}_{u}(t)=0$ for $u<t$):
$F_{t}^{i}(u,t)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}_{t}(u),$
@@
$\partial_{t}F_{t}^{i}(u,s)=\widetilde{X}^{i}_{t}(u)^{\top}P^{i}(t)\,\widetilde{X}^{i}_{t}(s),$
@@
The formula (A.19) is explicit in the gain kernel $\widetilde{X}^{i}$: once $\widetilde{X}^{i}$ is computed from (A.12), the filtering kernel $F^{i}$ requires no further fixed-point iteration. In the numerical scheme of Section 4, the integral $\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr$ is accumulated forward alongside the $\widetilde{X}^{i}$-computation, and the two boundary terms are set at each new observation time.